OpenStudy (anonymous):

In a circle, a 90 degree sector has area of 64pi ftsqd. what is the radius of the circle?

OpenStudy (anonymous):

if a circle has 360 degrees, 90 degrees is one forth of it

OpenStudy (anonymous):

so the total area is 4 times 64pi or 256 pi

OpenStudy (anonymous):

A=pi times r^2 256pi=(pi)r^2 256=r^2 r=16